PS1 IPP Czar Logs for the week YYYY.MM.DD - YYYY.MM.DD

(Up to PS1 IPP Czar Logs)

Monday : 2016-03-07

  • 17:00 CZW: Restarted ipp pantasks servers. Still waiting for Haydn to reboot ipp103 before setting them to run (or 7 pm).
  • 17:37 CZW: Processing back on.
  • 19:55 EAM: restarted gmond on ipp103

Tuesday : 2016-03-08

  • 9:54 HAF: Serge/Mops reports that ps_33_2316 is not in the correct order - there are 5 exposures, 2 visit 1's which I suspect cause problems. I requeued:
    difftool -dbname gpc1 -definewarpwarp -exp_id 1057357 -template_exp_id 1057374 -backwards -set_workdir neb://@HOST@.0/gpc1/OSS.nt/2016/03/08 -set_dist_group SweetSpot -set_label OSS.nightlyscience -set_data_group OSS.20160308 -set_reduction SWEETSPOT -simple -rerun
    difftool -dbname gpc1 -definewarpwarp -exp_id 1057391 -template_exp_id 1057409 -backwards -set_workdir neb://@HOST@.0/gpc1/OSS.nt/2016/03/08 -set_dist_group SweetSpot -set_label OSS.nightlyscience -set_data_group OSS.20160308 -set_reduction SWEETSPOT -simple -rerun
    
    
  • 14:00 CZW: Started ipplanl/pv3update pantasks to do last set of updates for STSCI. This is only a few hundred exposures worth of data, so it should be finished by the end of the day.
  • 18:50 HAF: cleared up 7 diff faults, they all have data error code 36b6 - set to 42 quality.

Wednesday : YYYY.MM.DD

  • 7:45 HAF: serge reports no warps have been done. IT seems a very high number of things faulted (looks like, due to nebulous, type 2, like 65% of chips), I am restarting pantasks. will investigate after that

Thursday : 2016.03.10

  • 18:10 EAM: restarting pantasks for the night

Friday : 2016.03.11

  • 21:00 EAM: restarting pantasks for the night

Saturday : YYYY.MM.DD

  • 20:00 EAM: restarting pantasks for the night

Sunday : YYYY.MM.DD

  • 20:40 EAM: restarting pantasks for the night